OpenStudy (anonymous):

How many ions are present in 30.0mL of 0.600M Na2CO3 solution

OpenStudy (anonymous):

Molarity = moles / volume in dm3 or litre 1st of all convert volume 30 mL into Litre or dm3 by dividing it with 1000. then it becomes 30/1000 = 0.03 Now find the moles by arranging the above equation, i.e moles = molarity * volume moles = 0.6 * 0.03 moles = 0.018 now we have got the moles of Na2CO3 1 mole of Na2CO3 = 3 moles of total ions. (2NA+ and CO-2) 0.018 moles of Na2CO3 = 3*0.018 0.018 moles of Na2CO3 = 0.054 moles of ions now apply the formula, i.e number of atoms or ions = moles * Na (avogadros number.) number of ions = moles of ions * Na (avogadro's number) number of ions = 0.054 * 6.022 * 10^{3} number of ions = 3.025 * 10^{22}

OpenStudy (aaronq):

good work, but next time dont solve it. Just tell them the steps they need to take.

OpenStudy (anonymous):

ok and thanks